OpenStudy (anonymous):

Help, medal will be awarded! Line BC has an equation of a line y = 2x + 3, and line EF has an equation of a line y = x + 4. These two equations represent parallel lines because the slopes of the lines are equal perpendicular lines because the slopes are opposite reciprocals the same line when graphed lines that intersect, but not at a right angle

OpenStudy (anonymous):

Lines that intersect, but not at a right angle

OpenStudy (anonymous):

The slopes are neither the same nor are they opposite reciprocals, which eliminates the first three answers

OpenStudy (bradely):

y = 2x+3,y = x+4 Observe the slopes m1=2,m2=1 They are not equal so they are not parallel m1*m2 not equal to -1 So they are intersect. Option 4 correct Source: http://www.mathskey.com/question2answer/